OpenStudy (anonymous):

2 square root of x-4 +16= 32

OpenStudy (anonymous):

can u put parenthesis around what's under the square root symbol?

OpenStudy (anonymous):

I'm guessing you meant 2 sqr root(x-4)+16=32 first divide each side by 2 yielding: sqr root(x-4) +8=16 subtract 8 from each side: sqr root(x-4)=8 Square both sides: x-4=64 add 4 to each side x=68